### v1

@@ Abstract
Visual navigation in indoor environments is a topic within the field of Embodied AI research which focuses on an agent/robot’s ability to follow instructions to navigate within unknown environments towards a goal. Approaching this problem typically requires an agent/robot to: 1) understand and remember the environment its been placed in; 2) interpret natural language instructions; and 3) use information from the formers to decide on a series of actions to adequately follow the instructions given to it [39].
@@
While the premise of this task is straightforward, different variants have been introduced over the years, which can be broadly classified based on communication complexity (single/multi-turn interaction), task objective (action/goal-directed), and action space (discrete/continuous spaces) [39, 10, 4, 27].
@@
Within this framework, Visual Language Navigation (VLN) is generally a single-turn, action-directed task, with discrete or continuous action spaces depending on the task variant [4, 16]. VLN in continuous indoor environments (VLN-CE) has gained significant attention recently due to its alignment with increasingly probable real-world applications, such as home robotic assistants [16]. However, the commonly used task datasets and simulators for existing VLN tasks are largely static and lack the dynamic complex features in real-world scenarios, such as moving obstacles and changing spaces. In real settings, humans and other entities often move within the same space that a robot is navigating, which require agents to not only follow instructions but also predict the future positions of these dynamic obstacles and adjust their routes accordingly at inference time. These are capabilities essential for success in other related navigation tasks like SOON, HANNA, VLNA, and VDN. [40, 37, 20, 25, 30].
@@
To narrow this gap, we introduce Adaptive Visual Language Navigation (AdaVLN), a task extension of the VLN-CE problem, to incorporate moving human obstacles into the widely-used Habitat-Matterport3D environments [26]. Additionally, we propose a ”freeze-time” mechanism for simulations in AdaVLN, where the otherwise constantly-running simulation is paused while the agent is processing decisions, ensuring fair comparisons across varying hardware speeds.
@@
Specifically, we introduce the following two tools to enable research in this topic: AdaSimulator: A simulator offering physics-based 3D environments with dynamically moving obstacles like humans and accurate mobile robot movements (refer to Figure 1 and 2). AdaSimulator is based on IsaacSim [19] and is built to be compatible with Matterport3D environments [6] and supports easy customisation of human spawn points, and pathing logic. AdaR2R: A sample variant of the R2R [4] and Matterport3D datasets that includes spawn points and trajectories for dynamic obstacles, adding another layer of realism to the navigation task.
@@ 1.1 Visual Language Navigation
Over the years, the VLN task has evolved and produced several variants, generally aiming to narrow the gap between simulated environments and the real-world scenarios that an actual robot might encounter. The original Visual Language Navigation task and Room-to-Room (R2R) dataset were introduced by Anderson et al. [4] and requires robots to navigate in static 3D environments given a single initial instruction. At every navigation step, the robot is provided with a 360-degree panoramic RGB-D view of its surroundings, and has to choose from pre-determined neighbouring nodes to teleport to. It popularised the use of the Matterport3D scan dataset [6] as a source of realistic 3D environments and provided the original Matterport3D Simulator. This simulator was later adapted to similar tasks in static environments, such as Scenario-Oriented Object Navigation (SOON) [40] and Remote Embodied Visual Referring Expressions (REVERIE) [25].
@@
Soon after, expansions to the original R2R datasets, e.g. R4R [14], RxR [17], were created to diversify and increase the difficulty of the navigation tasks. In parallel, new tasks emerged that shifted the focus to different complexities and problems within the field of visual navigation. The Embodied Question Answering (EQA) task [37] and Vision-and-Language Navigation with Actions (VLNA) [20] were introduced as related tasks in which agents were challenged not only to navigate but also to answer questions or perform actions based on the visual scene.
@@
The Habitat Sim simulator [24, 29, 28] and Habitat-Matterport3D [26] mesh datasets were later introduced, which provided a framework for conducting experiments in full physics-enabled 3D environments. Krantz el al [16] integrated this to extend the VLN task to continuous action spaces (VLN-CE), where robots had to navigate by making ’low-level’ movement decisions (turn left/right 15 degrees, move forward 0.25m, stop). RxR-Habitat competitions focusing on the VLN-CE task have since been organised in multiple years of the CVPR conference, which implement slight variants to the task parameters [9, 2]. This task further closed the sim2real gap, and has led to research into new world state modelling techniques [1, 33, 36] and long-term navigation planning ideas [35, 32] specific to it. [12] also showed the natural link between the discrete and continuous variants, and how they can complement each other when solving either task.
@@
Recent work by Li et al. [18] introduced the Human-Aware MP3D (HA3D) simulator for discrete action spaces, with perspective-specific animations of humans in Matterport3D environments and a corresponding Human-Aware R2R dataset, requiring robots to parse these dynamic elements as part of their navigation instructions. The paper also introduced the inclusion of collision statistics into the metrics of VLN experiments like success rate etc.
@@ 1.2 Collision Avoidance during Navigation
Collision avoidance in robotics is a well-researched topic, where the idea of local/offline path-planning is especially important since robots are expected to perform well in unknown environments. This typically requires the model to predict how their world state will change along a trajectory in order to preemptively avoid obstacles when planning their path. Traditional methods in robotics research for choosing safe paths used Velocity Obstacles [8] to calculate potential collision paths based on both the movement of surrounding objects and the robot itself. Reciprocal Velocity Obstacles extended this to multi-agent scenarios common in real-world scenarios. Methods that extract motion information from RGB-D data was also demonstrated in [11]. Newer prediction methods also made use of reinforcement learning techniques, models dynamic obstacles as variable-sized ellipsoids, and considers agent-human interaction dynamics to ground path predictions [31, 23, 5]. These were typically combined with grid-based [7], graph-based [34], or 3D-based methods [13] of modelling the world on-the-fly provided these robots with rich representations of its surroundings for better motion prediction and environment understanding.
Visual Language Navigation has since integrated many of the above ideas for future planning and obstacle prediction. Notably, DREAMWALKER [32] introduced the use of mental simulations in order to predict the environment along candidate trajectories. [1] focused on obstacle avoidance using dynamic topological planning, and later on Jeong et al. [15] introduced the VLN-CM agent, which leverages depth maps to predict expected occupancy maps along their candidate trajectories.
@@ 2.2 Observations/Actions of Robots
At each navigation step , the robot observes an egocentric 115-degree front-facing view of its surroundings in the form of an RGB-D image [18], as seen in Figure 3. Based on this observation and its current state, the robot can choose from one of four possible actions:
@@ 3.1 AdaSimulator
All simulation scenarios use a two-wheeled NVIDIA Jetbot, controlled via differential controllers for physics-based movement. All egocentric observations are rendered through IsaacSim’s Replicator Core, using the Jetbot’s attached camera for render perspective. Dynamic human obstacles are introduced into the environment via a customized version of the omni.anim.people [21] extension. A ROS2 interface is also provided, allowing agents to extract RGB-D observations from the simulator and send control commands.
@@ 3.2 AdaR2R (Sample)
AdaR2R (Sample) is an example dataset containing 9 navigation episodes across 3 HM3Dv2 example scenes [26]. Snapshots of these episode’s environments and human obstacles are shown in Figure 4. It modifies the original R2R dataset format to include configurations for human spawn points, path waypoints, and movement parameters. The example configurations have been manually set up to include 1-2 humans per episode, and their paths waypoints are chosen such that they will directly interfere with straight-line paths between critical nodes provided in the reference path. However, these interferences are never permanent, and there will either always be an alternative route that curves around the obstacle, or the obstacle will eventually move away as part of its patrol.
@@ 4.1 Evaluation Protocol
To give a baseline demonstration of the task and use of the simulator, we evaluate a baseline agent’s ability to navigate to its goal and avoid collision with both humans and environmental obstacles. Established evaluation metrics for VLN tasks typically focus on the navigation performance of agents [3, 4, 38]. Due to our focus on introducing a new simulation framework rather then an agent, we will instead look at our baseline agent’s collisions with environmental and human obstacles instead. Navigation Collisions (NC) records the ratio of the total amount of time an agent is in collision with either a human or static environmental obstacles (walls, furnitures etc.) to the total navigation time. We also break it down into Human Navigation Collisions (HNC) and Environmental Navigation Collisions. We will also do a qualitative analysis of our baseline agent’s observations and actions for several navigation episodes.
@@ 4.3 Baseline GPT Agent
We test a simple agent that uses the GPT-4o-mini [22] multi-modal foundational model. At each navigation step, the robot was told to generate semantic observations from the images, before generating a long-term plan to progress in the task. The robot then looks back on previous steps, then reasons about the logical next step required to continue along the plan, before finally making a decision. Predictions were limited to 200 output tokens per navigation step, and only RGB images used as input (after down-scaling from the render product’s 1280x720 to 640x360 pixels) to reduce complexity and token usage. Implementation details are available in the project repository.
@@ 5 Conclusion and Future Work
r*3c
Episode Environmental NC Human NC Combined NC
1 0.77 0.01 0.78
2 0.69 0.01 0.70
3 0.91 0.01 0.91
4 0.93 0.00 0.93
5 0.86 0.00 0.86
6 0.76 0.00 0.76
7 0.00 0.00 0.00
8 0.71 0.08 0.78
9 0.00 0.00 0.00
Average 0.63 0.01 0.64

### v2

@@ Abstract
Visual navigation in indoor environments is a topic within the field of Embodied AI research which focuses on an agent/robot’s ability to follow instructions to navigate within unknown environments towards a goal. Approaching this problem typically requires an agent/robot to: 1) understand and remember the environment its been placed in; 2) interpret natural language instructions; and 3) use information from the formers to decide on a series of actions to adequately follow the instructions given to it [Zhang_Ma_Li_Qiao_Wang_Chai_Wu_Bansal_Kordjamshidi_2024].
@@
While the premise of this task is straightforward, different variants have been introduced over the years, which can be broadly classified based on communication complexity (single/multi-turn interaction), task objective (action/goal-directed), and action space (discrete/continuous spaces) [Zhang_Ma_Li_Qiao_Wang_Chai_Wu_Bansal_Kordjamshidi_2024, Gu_Stefani_Wu_Thomason_Wang_2022, Anderson_Wu_Teney_Bruce_Johnson_Sünderhauf_Reid_Gould_Hengel_2018, Raychaudhuri_Ta_Ashton_Chang_Wang_Bucher_2024].
@@
Within this framework, Visual Language Navigation (VLN) is generally a single-turn, action-directed task, with discrete or continuous action spaces depending on the task variant [Anderson_Wu_Teney_Bruce_Johnson_Sünderhauf_Reid_Gould_Hengel_2018, Krantz_Wijmans_Majumdar_Batra_Lee_2020]. VLN in continuous indoor environments (VLN-CE) has gained significant attention recently due to its alignment with increasingly probable real-world applications, such as home robotic assistants [Krantz_Wijmans_Majumdar_Batra_Lee_2020]. However, the commonly used task datasets and simulators for existing VLN tasks are largely static and lack the dynamic complex features in real-world scenarios, such as moving obstacles and changing spaces. In real settings, humans and other entities often move within the same space that a robot is navigating, which require agents to not only follow instructions but also predict the future positions of these dynamic obstacles and adjust their routes accordingly at inference time. These are capabilities essential for success in other related navigation tasks like SOON, HANNA, VLNA, and VDN. [Zhu_Liang_Zhu_Chang_Liang_2021, Wijmans_Datta_Maksymets_Das_Gkioxari_Lee_Essa_Parikh_Batra_2019, Nguyen_Dey_Brockett_Dolan_2019, Qi_Wu_Anderson_Wang_Wang_Shen_Hengel_2020, Thomason_Murray_Cakmak_Zettlemoyer_2019].
@@
To narrow this gap, we introduce Adaptive Visual Language Navigation (AdaVLN), a task extension of the VLN-CE problem, to incorporate moving human obstacles into the widely-used Habitat-Matterport3D environments [Ramakrishnan_Gokaslan_Wijmans_Maksymets_Clegg_Turner_Undersander_Galuba_Westbury_Chang_2021]. Additionally, we propose a ”freeze-time” mechanism for simulations in AdaVLN, where the otherwise constantly-running simulation is paused while the agent is processing decisions, ensuring fair comparisons across varying hardware speeds.
@@
Specifically, we introduce the following two tools to enable research in this topic: AdaSimulator: A simulator offering physics-based 3D environments with dynamically moving obstacles like humans and accurate mobile robot movements (refer to Figure 1 and 2). AdaSimulator is based on IsaacSim [makoviychuk2021isaacgymhighperformance] and is built to be compatible with Matterport3D environments [Chang_Dai_Funkhouser_Halber_Nießner_Savva_Song_Zeng_Zhang_2017] and supports easy customisation of human spawn points, and pathing logic. AdaR2R: A sample variant of the R2R [Anderson_Wu_Teney_Bruce_Johnson_Sünderhauf_Reid_Gould_Hengel_2018] and Matterport3D datasets that includes spawn points and trajectories for dynamic obstacles, adding another layer of realism to the navigation task.
@@ 1.1 Visual Language Navigation
Over the years, the VLN task has evolved and produced several variants, generally aiming to narrow the gap between simulated environments and the real-world scenarios that an actual robot might encounter. The original Visual Language Navigation task and Room-to-Room (R2R) dataset were introduced by Anderson et al. [Anderson_Wu_Teney_Bruce_Johnson_Sünderhauf_Reid_Gould_Hengel_2018] and requires robots to navigate in static 3D environments given a single initial instruction. At every navigation step, the robot is provided with a 360-degree panoramic RGB-D view of its surroundings, and has to choose from pre-determined neighbouring nodes to teleport to. It popularised the use of the Matterport3D scan dataset [Chang_Dai_Funkhouser_Halber_Nießner_Savva_Song_Zeng_Zhang_2017] as a source of realistic 3D environments and provided the original Matterport3D Simulator. This simulator was later adapted to similar tasks in static environments, such as Scenario-Oriented Object Navigation (SOON) [Zhu_Liang_Zhu_Chang_Liang_2021] and Remote Embodied Visual Referring Expressions (REVERIE) [Qi_Wu_Anderson_Wang_Wang_Shen_Hengel_2020].
@@
Soon after, expansions to the original R2R datasets, e.g. R4R [Jain_Magalhaes_Ku_Vaswani_Ie_Baldridge_2019], RxR [Ku_Anderson_Patel_Ie_Baldridge_2020], were created to diversify and increase the difficulty of the navigation tasks. In parallel, new tasks emerged that shifted the focus to different complexities and problems within the field of visual navigation. The Embodied Question Answering (EQA) task [Wijmans_Datta_Maksymets_Das_Gkioxari_Lee_Essa_Parikh_Batra_2019] and Vision-and-Language Navigation with Actions (VLNA) [Nguyen_Dey_Brockett_Dolan_2019] were introduced as related tasks in which agents were challenged not only to navigate but also to answer questions or perform actions based on the visual scene.
@@
The Habitat Sim simulator [puig2023habitat3, szot2021habitat, habitat19iccv] and Habitat-Matterport3D [Ramakrishnan_Gokaslan_Wijmans_Maksymets_Clegg_Turner_Undersander_Galuba_Westbury_Chang_2021] mesh datasets were later introduced, which provided a framework for conducting experiments in full physics-enabled 3D environments. Krantz el al [Krantz_Wijmans_Majumdar_Batra_Lee_2020] integrated this to extend the VLN task to continuous action spaces (VLN-CE), where robots had to navigate by making ’low-level’ movement decisions (turn left/right 15 degrees, move forward 0.25m, stop). RxR-Habitat competitions focusing on the VLN-CE task have since been organised in multiple years of the CVPR conference, which implement slight variants to the task parameters [habitatrxrcompetition, an20221stplacesolutionsrxrhabitat]. This task further closed the sim2real gap, and has led to research into new world state modelling techniques [an2024etpnavevolvingtopologicalplanning, wang2023graphbasedenvironmentrepresentation, wang2023gridmmgridmemorymap] and long-term navigation planning ideas [wang2024lookaheadexplorationneuralradiance, Wang_Liang_Van_Gool_Wang_2023] specific to it. [hong2022bridginggaplearningdiscrete] also showed the natural link between the discrete and continuous variants, and how they can complement each other when solving either task.
@@
Recent work by Li et al. [Li_Li_Cheng_Dong_Zhou_He_Dai_Mitamura_Hauptmann_2024] introduced the Human-Aware MP3D (HA3D) simulator for discrete action spaces, with perspective-specific animations of humans in Matterport3D environments and a corresponding Human-Aware R2R dataset, requiring robots to parse these dynamic elements as part of their navigation instructions. The paper also introduced the inclusion of collision statistics into the metrics of VLN experiments like success rate etc.
@@ 1.2 Collision Avoidance during Navigation
Collision avoidance in robotics is a well-researched topic, where the idea of local/offline path-planning is especially important since robots are expected to perform well in unknown environments. This typically requires the model to predict how their world state will change along a trajectory in order to preemptively avoid obstacles when planning their path. Traditional methods in robotics research for choosing safe paths used Velocity Obstacles [Fiorini_Shiller_1998] to calculate potential collision paths based on both the movement of surrounding objects and the robot itself. Reciprocal Velocity Obstacles extended this to multi-agent scenarios common in real-world scenarios. Methods that extract motion information from RGB-D data was also demonstrated in [Herbst_Ren_Fox_2013]. Newer prediction methods also made use of reinforcement learning techniques, models dynamic obstacles as variable-sized ellipsoids, and considers agent-human interaction dynamics to ground path predictions [Truong_Ngo_2017, Pfeiffer_Shukla_Turchetta_Cadena_Krause_Siegwart_Nieto_2018, Castillo-Lopez_Sajadi-Alamdari_Sanchez-Lopez_Olivares-Mendez_Voos_2018]. These were typically combined with grid-based [Elfes_2013], graph-based [Wang_Wu_Yao_Wang_2023], or 3D-based methods [Hornung_Wurm_Bennewitz_Stachniss_Burgard_2013] of modelling the world on-the-fly provided these robots with rich representations of its surroundings for better motion prediction and environment understanding.
Visual Language Navigation has since integrated many of the above ideas for future planning and obstacle prediction. Notably, DREAMWALKER [Wang_Liang_Van_Gool_Wang_2023] introduced the use of mental simulations in order to predict the environment along candidate trajectories. [an2024etpnavevolvingtopologicalplanning] focused on obstacle avoidance using dynamic topological planning, and later on Jeong et al. [Jeong_Kang_Kim_Zhang_2024] introduced the VLN-CM agent, which leverages depth maps to predict expected occupancy maps along their candidate trajectories.
@@ 2.2 Observations/Actions of Robots
At each navigation step , the robot observes an egocentric 115-degree front-facing view of its surroundings in the form of an RGB-D image [Li_Li_Cheng_Dong_Zhou_He_Dai_Mitamura_Hauptmann_2024], as seen in Figure 3. Based on this observation and its current state, the robot can choose from one of four possible actions:
@@ 3.1 AdaSimulator
All simulation scenarios use a two-wheeled NVIDIA Jetbot, controlled via differential controllers for physics-based movement. All egocentric observations are rendered through IsaacSim’s Replicator Core, using the Jetbot’s attached camera for render perspective. Dynamic human obstacles are introduced into the environment via a customized version of the omni.anim.people [nvidia2024omnianimpeople] extension. A ROS2 interface is also provided, allowing agents to extract RGB-D observations from the simulator and send control commands.
@@ 3.2 AdaR2R (Sample)
AdaR2R (Sample) is an example dataset containing 9 navigation episodes across 3 HM3Dv2 example scenes [Ramakrishnan_Gokaslan_Wijmans_Maksymets_Clegg_Turner_Undersander_Galuba_Westbury_Chang_2021]. Snapshots of these episode’s environments and human obstacles are shown in Figure 4. It modifies the original R2R dataset format to include configurations for human spawn points, path waypoints, and movement parameters. The example configurations have been manually set up to include 1-2 humans per episode, and their paths waypoints are chosen such that they will directly interfere with straight-line paths between critical nodes provided in the reference path. However, these interferences are never permanent, and there will either always be an alternative route that curves around the obstacle, or the obstacle will eventually move away as part of its patrol.
@@ 4.1 Evaluation Protocol
To give a baseline demonstration of the task and use of the simulator, we evaluate a baseline agent’s ability to navigate to its goal and avoid collision with both humans and environmental obstacles. Established evaluation metrics for VLN tasks typically focus on the navigation performance of agents [Anderson_Chang_Chaplot_Dosovitskiy_Gupta_Koltun_Kosecka_Malik_Mottaghi_Savva_et_al_2018, Anderson_Wu_Teney_Bruce_Johnson_Sünderhauf_Reid_Gould_Hengel_2018, Yue_Zhou_Xie_Zhang_Yan_Yin_2024]. Due to our focus on introducing a new simulation framework rather then an agent, we will instead look at our baseline agent’s collisions with environmental and human obstacles instead. Navigation Collisions (NC) records the ratio of the total amount of time an agent is in collision with either a human or static environmental obstacles (walls, furnitures etc.) to the total navigation time. We also break it down into Human Navigation Collisions (HNC) and Environmental Navigation Collisions. We will also do a qualitative analysis of our baseline agent’s observations and actions for several navigation episodes.
@@ 4.3 Baseline GPT Agent
We test a simple agent that uses the GPT-4o-mini [openai2024gpt4omini] multi-modal foundational model. At each navigation step, the robot was told to generate semantic observations from the images, before generating a long-term plan to progress in the task. The robot then looks back on previous steps, then reasons about the logical next step required to continue along the plan, before finally making a decision. Predictions were limited to 200 output tokens per navigation step, and only RGB images used as input (after down-scaling from the render product’s 1280x720 to 640x360 pixels) to reduce complexity and token usage. Implementation details are available in the project repository.
@@ 5 Conclusion and Future Work
| Episode | Environmental NC | Human NC | Combined NC |
| 1 | 0.77 | 0.01 | 0.78 |
| 2 | 0.69 | 0.01 | 0.70 |
| 3 | 0.91 | 0.01 | 0.91 |
| 4 | 0.93 | 0.00 | 0.93 |
| 5 | 0.86 | 0.00 | 0.86 |
| 6 | 0.76 | 0.00 | 0.76 |
| 7 | 0.00 | 0.00 | 0.00 |
| 8 | 0.71 | 0.08 | 0.78 |
| 9 | 0.00 | 0.00 | 0.00 |
| Average | 0.63 | 0.01 | 0.64 |
6 Acknowledgements
This project is supported via the NVIDIA Academic Grant, including the A100 80GB GPU hardwares, NVIDIA IsaacSim software and Saturn Enterprise cloud.